# Investigation of the Challenges of Underwater-Visual-Monocular-SLAM

### v1

@@ 4.2 AUV Datasets
The AUVs have a circular-arranged active lighting system, comprised of 8 LED-compounds cast in rasin (see Fig. 1) [Sticklus et al., 2017, Song et al., 2021a]. Specifically, the AUVs are equipped with a dome port camera, which was again calibrated in air with a fisheye model. Subsequently, it was centered [She et al., 2019, She et al., 2022] to avoid refraction effects. We then overtake the underwater fisheye parameters estimated by Metashapes Photoscan into an adapted Colmap [Schönberger et al., 2016, Schönberger and Frahm, 2016] version to establish ground truth with an offline reconstruction method. In the latter process, the navigation data is fused into the visual reconstruction using a prose-graph-approach. For evaluation, the images are then undistorted and provided in canonical pinhole space for the SLAM algorithms. Finally, the results from Colmap’s sparse reconstruction serve as the ground truth poses.
@@ 5 Evaluation
We categorized failures into three types: not initializing (NOT INIT), initializing but losing track (TR-Lost), and complete failure (FAILED). Not initializing means the method could not start tracking the camera pose. Initializing but losing track indicates that the method began tracking but eventually lost the camera pose without recovery. If the camera poses are lost but the map has enough keyframes, the algorithm is considered successful. Complete failure means the method provided no output (e.g., LSD-SLAM). BADSLAM may fail to start if the estimated depth map is inaccurate. We used a solid alignment approach to match up the SLAM estimates and the ground truth. We initially temporally aligned the SLAM estimates and the ground truth interpolating the latter such that for each ground truth pose there is a corresponding estimated pose. Afterwards, we employed the $\mathbb{SIM}3$ [Allen-Blanchette et al., 2014] Umeyama [Umeyama, 1991] alignment technique, which considers scale, translation, and rotation, to accurately align the result trajectory with the ground truth in space. Specifically, we optimize
@@ 6 Results
According to the findings presented in Annex A, the experiments conducted on the A1, A2, and A3 datasets revealed that the algorithms encountered challenges and exhibited poor performance due to various factors. Specifically, for the A1 dataset, one of the main issues identified was the lack of sufficient overlap between frames. This insufficient overlap hindered the algorithms’ ability to establish robust correspondences and accurately estimate the robot’s trajectory. On the other hand, for the A2 dataset, although there was good overlap between frames, the presence of unfavorable water and light conditions posed significant difficulties for the algorithms. These conditions, such as poor visibility, light scattering, and limited lighting, adversely affected the algorithms’ ability to accurately estimate depth and track the robot’s movement. Furthermore as for the case of the A3 dataset, the challenges were further compounded by the presence of low texture areas in addition to the water and lighting conditions. Low texture areas, which lack distinctive visual features, make it challenging for SLAM algorithms to establish reliable correspondences and accurately estimate the robot’s trajectory in those regions. In fact, while the initialization trajectory of A3 helped for the initialization of the ORB-SLAMS and the median method correction was the most qualitative one for the A2 data-set, the ORB-SLAMS lost the track in the majority of the frames.
The experiments on the T8-13 dataset, which consisted of a tank with water, have shown that light conditions are critical for successful SLAM performance. The light cones produced from the artificial lights, with and without the sunlight, had a significant impact on the images, resulting in failure of the SLAM methods. This is due to the fact that the presence of water causes light to being refracted, attenuated and scattered. While we controlled for the refraction effects, by centering the camera in the dome, the two latter effects lead to changes in image appearance and the degradation of image quality. In addition, the absorption and scattering of light in water varies depending on the wavelength and scene depth, which can affect the accuracy of visual odometry and feature tracking.
@@
The BADSLAM and GRADSLAM algorithms did not work with either the UDepth or UW-Net depth estimators in underwater scenarios and with Monodepth2 for the in-air sets. To ascertain whether the depth estimators were the root of the issue, we used Colmap for depth estimation and discovered that BADSLAM was capable of at least initializing itself in the A1/2 and A3 datasets and in the T in-air sets it succeeded alongside GRADSLAM. In regard to the UW-Net, U-Depth, and Monodepth2 depth estimators, it is important to note that they are primarily designed for forward-looking camera settings and not specifically for top-down views. Unsurprisingly, they do not perform optimally when applied to top-down views, such as those encountered in underwater environments (see Fig. 5).
@@ 7 Conclusion
In this paper, we conducted an investigation of the challenges of underwater monocular visual SLAM. To this end, we prepared several AUV-based and controlled lab-datasets. All geometric distortions are controlled for in those sets, while they are all taken in extremely low visibility and harsh light conditions. This allows for an in-depth investigation of the impact of radiometric distortions in the underwater setting. The ground truth in the lab sets is established with a custom-build external reference system, while the AUV sets – lacking external reference – are offline-reconstructed with a modified version Colmap. First, we showed that all selected SLAM algorithms successfully run on in-air tank-datets with good performances. Then, we evaluated several combinations of SLAM systems and preprocessing methods on all datasets. We found that no SLAM system is able to complete the real AUV-Datasets. Also from the tank datasets, only the homogeneously illuminated Sun-settings could be completed. Here, we found that the pre-processing approaches showed some initial improvements of the SLAM performance in the visually adversarial underwater environments. In addition, we can also report mild improvements, when special initialization-maneuvers are carried out. The generalization of the pre-processing methods is a direction worth to further investigate, as they seem to be heavily tuned to certain assumptions / scenarios. Furthermore, providing depth-information dependent SLAM systems with corresponding top-down-view estimates also seems to be an interesting route. However we had to resort to depth maps established in an offline fashion, as the deep learning based estimators were tuned to different use cases. Hence, in the future, we will strive to preprocess underwater imagery to mitigate radiometric distortions and at the same time improve on underwater monocular depth estimation in order to leverage the already existing big potential of in-air SLAM approaches.

### v2

@@ 4.2 AUV Datasets
The AUVs have a circular-arranged active lighting system, comprised of 8 LED-compounds cast in rasin (see Fig. 1) [Sticklus et al., 2017, Song et al., 2021a]. Specifically, the AUVs are equipped with a dome port camera, which was again calibrated in air with a fisheye model. Subsequently, it was centered [She et al., 2019, She et al., 2022] to avoid refraction effects. We then overtake the underwater fisheye parameters estimated by Metashapes Photoscan into an adapted Colmap [Schönberger and Frahm, 2016] version to establish ground truth with an offline reconstruction method. In the latter process, the navigation data is fused into the visual reconstruction using a pose-graph-approach [She et al., 2023]. For evaluation, the images are then undistorted and provided in canonical pinhole space for the SLAM algorithms. Finally, the results from Colmap’s sparse reconstruction serve as the ground truth poses.
@@ 5 Evaluation
We categorized failures into three types: not initializing (NOT INIT), initializing but losing track (TR-Lost), and complete failure (FAILED). Not initializing means the method could not start tracking the camera pose. Initializing but losing track indicates that the method began tracking but eventually lost the camera pose without recovery. If the camera poses are lost but the map has enough keyframes, the algorithm is considered successful. Complete failure means the method provided no output (e.g., LSD-SLAM). BADSLAM may fail to start if the estimated depth map is inaccurate.
We used a solid alignment approach to match up the SLAM estimates and the ground truth. We initially temporally aligned the SLAM estimates and the ground truth interpolating the latter such that for each ground truth pose there is a corresponding estimated pose. Afterwards, we employed the $\mathbb{SIM}3$ [Allen-Blanchette et al., 2014] Umeyama [Umeyama, 1991] alignment technique, which considers scale, translation, and rotation, to accurately align the result trajectory with the ground truth in space. Specifically, we optimize
@@ 6 Results
According to the findings presented in Annex A, the experiments conducted on the A1, A2, and A3 datasets revealed that the algorithms encountered challenges and exhibited poor performance due to various factors. Specifically, for the A1 dataset, one of the main issues identified was the lack of sufficient overlap between frames. This insufficient overlap hindered the algorithms’ ability to establish robust correspondences and accurately estimate the robot’s trajectory. On the other hand, for the A2 dataset, although there was a good overlap between frames, the presence of unfavorable water and light conditions posed significant difficulties for the algorithms. These conditions, such as poor visibility, light scattering, and limited lighting, adversely affected the algorithms’ ability to accurately estimate depth and track the robot’s movement. Furthermore, as for the case of the A3 dataset, the challenges were further compounded by the presence of low texture areas in addition to the water and lighting conditions. Low texture areas, which lack distinctive visual features, make it challenging for SLAM algorithms to establish reliable correspondences and accurately estimate the robot’s trajectory in those regions. In fact, while the initialization trajectory of A3 helped for the initialization of the ORB-SLAMS and the median method performed the best on the A2 dataset, the ORB-SLAMS lost track in the majority of the frames.
The experiments on the T8-13 dataset, which consisted of a tank with water, have shown that light conditions are critical for successful SLAM performance. The light cones produced from the artificial lights, with and without the sunlight, had a significant impact on the images, resulting in failure of the SLAM methods. This is due to the fact that the presence of water causes light to being refracted, attenuated and scattered. While we controlled for the refraction effects, by centering the camera in the dome, the two latter effects led to changes in image appearance and the degradation of image quality. In addition, the absorption and scattering of light in water varies depending on the wavelength and scene depth, which can affect the accuracy of visual odometry and feature tracking.
@@
The BADSLAM and GRADSLAM algorithms did not work with either the UDepth or UW-Net depth estimators in underwater scenarios and with Monodepth2 for the in-air sets. To ascertain whether the depth estimators were the root of the issue, we used Colmap for depth estimation and discovered that BADSLAM was capable of at least initializing itself in the A1/2 and A3 datasets and in the T in-air sets it succeeded alongside GRADSLAM. In regard to the UW-Net, U-Depth, and Monodepth2 depth estimators, it is important to note that they are primarily designed for forward-looking camera settings and not specifically for top-down views. Unsurprisingly, they do not perform well when applied to top-down views, such as those encountered in underwater environments (see Fig. 5).
@@ 7 Conclusion
In this paper, we conducted an investigation of the challenges of underwater monocular visual SLAM. To this end, we prepared several AUV-based and controlled lab-datasets. All geometric distortions are controlled for in those sets, while they are all taken in extremely low visibility and harsh light conditions. This allows for an in-depth investigation of the impact of radiometric distortions in the underwater setting. The ground truth in the lab sets is established with a custom-build external reference system, while the AUV sets – lacking external reference – are offline-reconstructed with a modified version of Colmap. First, we showed that all selected SLAM algorithms successfully run on in air tank-dataets with good performances. Then, we evaluated several combinations of SLAM systems and preprocessing methods on all datasets. We found that no SLAM system is able to complete the real AUV-Datasets. Also from the tank datasets, only the homogeneously illuminated Sun-settings could be completed. Here, we found that the pre-processing approaches showed some initial improvements of the SLAM performance in the visually adversarial underwater environments. In addition, we can also report mild improvements, when special initialization-maneuvers are carried out. The generalization of the pre-processing methods is a direction worth to further investigate, as they seem to be heavily tuned to certain assumptions / scenarios. Furthermore, providing depth-information dependent SLAM systems with corresponding top-down-view estimates also seems to be an interesting route. However we had to resort to depth maps established in an offline fashion, as the deep learning based estimators were tuned to different use cases. Hence, in the future, we will strive to preprocess underwater imagery to mitigate radiometric distortions and at the same time improve on underwater monocular depth estimation in order to leverage the already existing big potential of in-air SLAM approaches.